A drinking glass is in the shape of a cylinder that is 60mm in diameter and 15cm tall. How many cubic centimeters of water will it hold?
Answer by stanbon(48522)

A drinking glass is in the shape of a cylinder that is 60mm in diameter and 15cm tall. How many cubic centimeters of water will it hold?
60mm = 6 cm
radius= (1/2)diameter
---------------
V=(base)(height)
V=(pi)(r^2)(height)
V=pi (3^2)15
V=135pi cubic centimeters
V is approx 424 cubic centimeters
Cheers,
Stan H.